9. When polyacrylonitrile is burned, toxic gases are released. Refering to Table 11.2 for the structur of acryloni trile, what toxic gas would you predict to be the product of combustion of this CH C (CN 10. What is the molar mass of polystyrene, if a molecule of polystyrene consists of 25,000 monomers? Addition Polymers er units n Alkene monomers Alkane polymer of n
Answer #1

(9)

The monomer of polyacryolonitrile is acrylonitrile CH2=CH-CN

When polyacrylonitrile is burned it gives a poisionous gas like HCN, hydrogen cyanide.

(10)

n Styrene ------------> Polystyrene

Molar mass of Styrene (C6H5CH=CH2) , a monomer = 104 g/mol

Molar mass of polystyrene having 25000 monomeric units = 25000 * 104 = 2.6 * 106 g/mol